OpenStudy (anonymous):

Use De Moivre's Thereom to compute (sqrt(3) + i)^4.

OpenStudy (jamesj):

So if \( z = \sqrt{3} + i \), what is \( z \) in polar form, \[ z = re^{i\theta} \] That is, what are \( r \) and \( \theta \).

OpenStudy (jamesj):

Once you know that, \[ z^4 = r^4 e^{4\theta i} \]

OpenStudy (anonymous):

Is it (r cos ø + r sin (ø)i ) ?

OpenStudy (anonymous):

in polar form?

OpenStudy (jamesj):

Right, same thing it turns out, \[ z = r( \cos \theta + i \sin \theta ) \] What value of r and theta gives your value of z?

OpenStudy (jamesj):

Then \( z^4 = r^4 ( \cos(4\theta) + i \sin(4\theta)) \).

OpenStudy (anonymous):

Ok, i think that \[r = \sqrt{a ^{2} + b ^{2}} = 2\]

OpenStudy (jamesj):

yes

OpenStudy (saifoo.khan):

whenever u are free @james http://openstudy.com/#/updates/4ed58b7be4b0bcd98ca6ebe8

OpenStudy (anonymous):

but i'm having trouble understanding how to find theta

OpenStudy (jamesj):

Draw a diagram and think about the tan of theta. I'm going to go to this other problem for a bit.

OpenStudy (anonymous):

ok :)

OpenStudy (jamesj):

found theta?

OpenStudy (anonymous):

i'm not sure, but i think its arctan of a/b, so arctan of \[1/\sqrt{3}\] ?

OpenStudy (anonymous):

so i think pi/6?

OpenStudy (jamesj):

Yes. And you should draw the diagram of z in the complex plane to convince yourself that's right. You should also see where that formula comes from. See it geometrically in your diagram. And then see it algebraically. If z = a + bi = r . cos(theta) + r. sin(theta) i Then b/a = sin(theta)/cos(theta) = tan(theta) => theta = arctan(b/a) For the record, you have to be quite careful with this formula. For example if z = 1 + i, then tan(theta) = 1. And if w = -1 - i, then tan(theta) also equals 1. But the theta are quite different.

OpenStudy (jamesj):

Anyway, for your z, tan(theta) = 1/sqrt(3). As theta lies in the first quadrant, it must be that theta = pi/6.

OpenStudy (anonymous):

OK, Thanks! I think I can handle it from there. :)